Telescope Operator: GENO Support Astronomrt: Michelle Observer: Zheng Cai, Fuyan Bian, Linhua Jiang, Ran Wang.

2:35: taking U_spec twilight flat. 0.5s exposure 30,000-15,000ADU

2:41 taking r and V twilight flat.

Collimate telescope:can not run the dofpia in idl, because there are so many the IDL_PATH setting in .cshrc comment out all of them.

Pointing the telescope: can not run the lbcrangebal. Michelle login lbto account to run the command.

04:15 go to the first U-V exposure. The seeing is about 1.0-1.3".

05:01 re-collimate the telescope and go to the first U-r exposure. The seeing is about 1.0-1.2"

05:36 re-collimate the telescope and go to the second U-V exposure. The seeing is about 0.9"

06:20 re-collimate the telescope and go to the second U-r exposure. The seeing is about 1.0"

06:59 re-collimate the telescope and go to the third U-V exposure

07:40 re-collimate the telescope and go to the third U-r exposure. The first two red images are bad. Some readout problems We stop the OB and take some bias. The redchip1 looks bad, and now is looks better. We took another set of biases, but the last red bias does not readout. Michelle has to reboot the LBC.

08:30 Restart the observing and go to the third U-r exposure

09:10 re-collimate the telescope and go to the fourth U-V exposure seeing is getting worse.

09:50 re-collimate the telescope and go to the fourth U-r exposure

10:33 re-collimate the telescope and go to the fifth U-V exposure, the sky become brighter and 3 out of 5 images are done.

-- FuyanBian - 22 May 2012
Topic revision: r3 - 22 May 2012, FuyanBian
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback